A farmhouse shelters 10 animals. Some are pigs some are ducks, Altogether there are 36 legs. How many of each animal are there?
Found 2 solutions by ewatrrr, ikleyn:
Answer by ewatrrr(24785)

 
Hi,
A farmhouse shelters 10 animals. Some are pigs some are ducks,
Altogether there are 36 legs. How many of each animal are there?
Let x be the number of pigs and (10-x) be the number of ducks.
4x + 2(10-x) = 36 legs
4x + 20 - 2x = 36
2x = 16
x = 8, the number of pigs, 2 is the number of ducks 10-8+
Checking our Answer:
4*8 + 2*2 = 36 checks.
Wish You the Best in your Studies.

Answer by ikleyn(52858)

.

a TWIN problem was solved for you 4 minutes ago at this link

https://www.algebra.com/algebra/homework/Systems-of-equations/Systems-of-equations.faq.question.1174011.html

https://www.algebra.com/algebra/homework/Systems-of-equations/Systems-of-equations.faq.question.1174011.html


Consider it as your TEMPLATE.


Having this template in front of you, solve the current problem in the same way.


E n j o y (!)

-----------------

To see many other similar solved problems, look into the lessons
    - Problem on two-wheel and three-wheel bicycles
    - Problem on animals at a farm
    - Problem on pills in containers
    - What type of problems are these?
in this site.